OpenStudy (anonymous):

HELP! I don't understand this at all :( To find out how long it will take for $500 to double when invested at 5% annual interest compounded twice a year, you solve the following equation: 1000 = 500 (1 + ) 2t. Use complete sentence to describe each step of math needed to solve this equation.

OpenStudy (anonymous):

what is in parenthesis?

OpenStudy (anonymous):

(1+ .05/2)

OpenStudy (anonymous):

1000 = 500 (1+(.05/2))^2t sorry i forgot the carrot!

OpenStudy (phi):

the formula is \[ 1000 = 500 (1+\frac{0.05}{2})^{2t} \]

OpenStudy (phi):

where t is in years.

OpenStudy (phi):

to solve, the first step is divide both sides by 500

OpenStudy (phi):

\[ 2= 1.025^{2t} \]

OpenStudy (anonymous):

how did you get 1.025^2t

OpenStudy (phi):

I divided 2 into 0.05 to get 0.025 and added it to 1

OpenStudy (phi):

one way to "bring down the exponent" 2t is use logarithms. Let's use log base 10: log(2)= log(1.025^2t) then use this very useful property \[ log(a^b)= b \cdot log(a) \]

OpenStudy (anonymous):

but i thought you were supposed to be dividing 500 into it

OpenStudy (phi):

we start with 1000= 500 (1+0.05/2)^(2t) divide both sides by 500 to get 2= (1+0.05/2)^(2t) divide 0.05 by 2: 2= (1+0.025)^(2t) add 1 + 0.025= 1.025 2= (1.025)^(2t) or simply 2= 1.025^(2t) follow?

OpenStudy (anonymous):

oh nevermind haha ignore what i jsut said

OpenStudy (anonymous):

i follow

OpenStudy (phi):

take the log (base 10 tho it does not matter which base we use) of both sides log(2)= log(1.025^(2t)) can you use the rule \[ log(a^b)= b \cdot log(a) \] to rewrite this equation?

OpenStudy (anonymous):

i dont know what base ten means... but i have no idea how to rewrite it like that

OpenStudy (anonymous):

would it be like log(2^10) = 10 * log(1.025^(2t))

OpenStudy (phi):

using the rule means match the rule to your expression. if you have log(2^10) you can rewrite it as 10*log(2) (2 matches a and 10 matches b in the rule). in this case match the rule to log(1.025^(2t))

OpenStudy (anonymous):

10*log(2) = 10*log(2)

OpenStudy (anonymous):

theyre the same

OpenStudy (phi):

in this case match the rule to log(1.025^(2t))

OpenStudy (phi):

forget 10*log(2) that was an example. look at log(1.025^(2t)) which part matches the a in the rule. which part matches the b in the rule

OpenStudy (anonymous):

the 1.025 part is a

OpenStudy (phi):

yes

OpenStudy (anonymous):

2t is b

OpenStudy (phi):

yes. so now use the rule to rewrite log(1.025^(2t))

OpenStudy (anonymous):

soooo 2t*log(1.025)

OpenStudy (phi):

yes. so back to the problem log(2)= log(1.025^(2t)) which is also (using the rule) log(2)= (2t)*log(1.025) now divide both sides by 2*log(1.025)

OpenStudy (phi):

like this \[ \frac{log(2)}{2\cdot log(1.025)}= \frac{t\cdot 2 \cdot log(1.025)}{2\cdot log(1.025)}\]

OpenStudy (phi):

\[ \frac{log(2)}{2\cdot log(1.025)}= \frac{t\cdot \cancel{2 \cdot log(1.025)}}{\cancel{2\cdot log(1.025)}} \]